OpenStudy (anonymous):

What is 8-5(1+ 6n)= -30

OpenStudy (anonymous):

Distribute the -5 to the ( ) to simplify...then solve for n

OpenStudy (anonymous):

how dont know to solve it. i need to know the steps.

OpenStudy (anonymous):

Step 1...distribute the -5 into the parenthesis. Gives this: 8 - 5 -30n = -30

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3 - 30n = -30 Combined the 8 - 5 to get 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3 - 30n = -30 3 - 3 - 30n = -30 - 3 Subtracted 3 from each side to get the variable alone -30n = -33 -30n/-30 = -33/-30 divide both sides by -30 to get n by itself n = 11/10

OpenStudy (anonymous):

does that make sense